OpenStudy (anonymous):

PLEASE I NEED HELP!whats the speed for the navy pier ferris wheel if it Takes 7-minute a revolution and the diameter is 140 feet & the height is 46m and the mass is 293kg

OpenStudy (anonymous):

i need help like big time all i know is that i think we might used D/T=v

OpenStudy (anikhalder):

1 rev in 7 mins Hence,in 1 min no of rev= 1/7 So, \[\omega(angular speed)=\frac{ 1 }{ 7}\times.105rad/\sec=0.015rad/\sec \] rad is radians not radius Now, radius of wheel=70 feet or 21.3 metres Therefore, vel. of a pt. on the circumference of the wheel= ang. speed x radius =0.015 x 21.3 =0.319 m/s